Shawn and his bike have a total mass of 47.5 kg. Shawn rides his bike 0.64 km in 12.9 min at a constant velocity.
The acceleration of gravity is 9.8 m/s2 . What is Shawn’s kinetic energy?
Answerinunitsof J
1
Expert's answer
2011-10-27T09:28:15-0400
Kinetic energy is described by the following formula
Ek=mv^2/2
where m - mass of an object (in our case it's mass of Shawn and his bike)
v - velocity of an object ( in our case it's velocity with which Shawn rides his bike)
velocity was constant during ride hence we can find it as v= S/T=0.64*1000/12.9*60=0.83 m/s
Hence we can substitute values into the formula
Ek=47.5*0.83^2/2=16.36 J